# A lower bound on the right-handed neutrino mass from wash-in leptogenesis

### v1

@@ Abstract
In closing, we remark that the RHN mass bounds in Eq. (27) are still subject to remaining theory uncertainties due to our rough treatment of the CPI. In our analysis, we assumed that no CPI will develop as long as the bound in Eq. (25) is satisfied, while baryon number overproduction via helicity decay at the electroweak scale is unavoidable if it is not satisfied. This treatment could be refined by explicitly including the effect of anomalous $q_{e}$ violation in the $q_{e}$ Boltzmann equation, in combination with a new transport equation for hypermagnetic helicity (e.g., along the lines of Appendix C in Ref. Domcke:2019mnd ). In general, a proper understanding of the onset of the CPI at large chiral charge $\mu_{Y,5}$, however, calls for dedicated analyses based on magnetohydrodynamical simulations of the chiral plasma at temperatures $T\sim T_{e}$. We leave such refinements for future work and conclude instead that wash-in leptogenesis provides an intriguing framework for the generation of the BAU that can successfully operate at RHN masses as low as $M_{N}\sim 7\,\textrm{TeV}$.
@@
Acknowledgments — The authors would like to thank Dietrich Bödeker, Kohei Kamada and Kyohei Mukaida for helpful discussions and comments. M. A. M. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG) Collaborative Research Centre “Neutrinos and Dark Matter in Astro- and Particle Physics” (SFB 1258). The work of K. S. and D. W. was supported by DFG through Research Training Group (Graduiertenkolleg) 2149: Strong and Weak Interactions — from Hadrons to Dark Matter.

### v2

@@ Abstract
In closing, we remark that the RHN mass bounds in Eq. (31) are still subject to remaining theory uncertainties due to our rough treatment of the CPI. In our analysis, we assumed that no CPI will develop as long as the bound in Eq. (29) is satisfied, while baryon number overproduction via helicity decay at the electroweak scale is unavoidable if it is not satisfied. This treatment could be refined by explicitly including the effect of anomalous $q_{e}$ violation in the $q_{e}$ Boltzmann equation, in combination with a new transport equation for hypermagnetic helicity (e.g., along the lines of Appendix C in Ref. Domcke:2019mnd). In general, a proper understanding of the onset of the CPI at large chiral charge $\mu_{Y,5}$, however, calls for dedicated analyses based on magnetohydrodynamical simulations of the chiral plasma at temperatures $T\sim T_{e}$. We leave such refinements for future work and conclude instead that wash-in leptogenesis provides an intriguing framework for the generation of the BAU that can successfully operate at RHN masses as low as $M_{N}\sim 7\,\textrm{TeV}$. Wash-in leptogenesis at such a low RHN mass scale requires the presence of a primordial charge asymmetry in right-handed electrons; the exploration of possible chargegenesis mechanisms capable of producing this primordial input charge deserves more attention in future work as well. Finally, it would be interesting to study extended scenarios that allow for the generation of nonzero lepton flavor asymmetries already at the stage of chargegenesis. Scenarios of this type, which would blur the line between the dynamical sectors responsible for lepton flavor and lepton number violation, promise to result in even lower RHN mass bounds and a smooth parametric interpolation between our results in this paper and RHN mass bounds in other leptogenesis scenarios. We encourage more work in this direction.
@@
Acknowledgments — The authors would like to thank Dietrich Bödeker, Kohei Kamada and Kyohei Mukaida for helpful discussions and comments. M. A. M. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG) Collaborative Research Centre “Neutrinos and Dark Matter in Astro- and Particle Physics” (SFB 1258).